# Conditionally valid Probabilistic Conformal Prediction

### v1

@@ Abstract
We develop a new method for creating prediction sets that combines the flexibility of conformal methods with an estimate of the conditional distribution $\textup{P}_{Y\mid X}$. Most existing methods, such as conformalized quantile regression and probabilistic conformal prediction, only offer marginal coverage guarantees. Our approach extends these methods to achieve conditional coverage, which is essential for many practical applications. While exact conditional guarantees are impossible without assumptions about the data distribution, we provide non-asymptotic bounds that explicitly depend on the quality of the available estimate of the conditional distribution. Our confidence sets are highly adaptive to the local structure of the data, making them particularly useful in high heteroskedasticity situations. We demonstrate the effectiveness of our approach through extensive simulations, showing that it outperforms existing methods in terms of conditional coverage and improves the reliability of statistical inference in a wide range of applications.
@@ 1 Introduction
Conformal predictions are commonly used to construct prediction sets. Under minimal assumptions, they offer finite-sample validity [40, 36]. However, significant challenges arise with high heteroskedasticity, often leading to incorrect inferences [11].
The split-conformal approach uses a set of $n$ calibration data points $\{(X_{k},Y_{k})\}_{k\in[n]}$ with $X_{k}\in\mathbb{R}^{d}$ and $Y_{k}\in\mathcal{Y}$ to create a prediction set $\mathcal{C}_{\alpha}(x)$ where $\alpha\in(0,1)$. For each $x\in\mathbb{R}^{d}$, the prediction set based on a conformity score function $V\colon\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$, is given by
@@
$\mathcal{C}_{\alpha}(x)=\left\{{y\in\mathcal{Y}\colon V(x,y)\leq Q_{1-\alpha}\left({\frac{1}{n+1}\sum\nolimits_{k=1}^{n}\delta_{V(X_{k},Y_{k})}+\frac{1}{n+1}\delta_{\infty}}\right)}\right\},$
@@
where $Q_{1-\alpha}$ represents $(1-\alpha)$-quantile of the adjusted empirical score distribution $\frac{1}{n+1}\sum_{k=1}^{n}\delta_{V(X_{k},Y_{k})}+\frac{1}{n+1}\delta_{\infty}$. If the calibrations data $\{(X_{k},Y_{k})\}_{k\in[n]}$ are drawn i.i.d. from a population distribution $\textup{P}_{X,Y}$, then for any new data point $(X_{n+1},Y_{n+1})\sim\textup{P}_{X,Y}$ sampled independently of the calibration data, the conformal theory ensures the marginal validity of $\mathcal{C}_{\alpha}(X_{n+1})$, meaning that
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)\geq 1-\alpha.$
Initially, most of the conformal methods focused on estimating a mean regression function for $Y\mid X$, to then construct a fixed-width band around it; see [39, 40]. However, as pointed out by [21], this marginal guarantee can hide significant discrepancies in the coverage of different regions of the input space $\mathbb{R}^{d}$. In particular, certain regions may be over-covered while others are under-covered. To solve this problem, it is necessary to construct adaptive prediction sets. Conditional conformal prediction methods allow the generation of confidence intervals that adapt to the test point under consideration $X_{n+1}$. For example, if a patient has certain characteristics, the goal is to guarantee that their treatment will be successful with a confidence of $1-\alpha$. Therefore, we want to adapt the guarantees to the profile of the individual instead of giving guarantees that only apply to the population as a whole. In particular, for any point $x\in\mathbb{R}^{d}$, the set $\mathcal{C}_{\alpha}(x)$ is described as conditional valid if
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x}\right)\geq 1-\alpha.$
Although conditional validity is a more desirable guarantee than marginal validity, it is difficult to achieve in practice without further assumptions about the data distribution. Indeed, it has been shown to be incompatible with the distribution-free setting [38, 25]. For practical purposes, however, an approximate conditional validity may be sufficient.
In regions with high heteroskedasticity, additional flexibility in the construction of prediction sets is useful. Methods based on quantile regression approaches were investigated in [30, 22]. These algorithms estimate the lower and upper conditional quantile regression functions $\hat{q}_{\alpha/2}$ and $\hat{q}_{1-\alpha/2}$; see e.g. [23, 2]. In [30], conditional confidence sets for $Y_{n+1}\mid X_{n+1}$ are constructed on the basis of the conformity assessment, which is determined by
$V(x,y)=\max\left\{{\hat{q}_{\alpha/2}(x)-y,y-\hat{q}_{1-\alpha/2}(x)}\right\}.$
Then, denoting by $\mu=\frac{1}{n}\sum_{k=1}^{n}\delta_{V(X_{k},Y_{k})}$; the prediction set is defined as
$\mathcal{C}_{\alpha}(x)=\left[{\hat{q}_{\alpha/2}(X)-Q_{(1-\alpha)(1+n^{-1})}(\mu),\hat{q}_{1-\alpha/2}(X)+Q_{(1-\alpha)(1+n^{-1})}(\mu)}\right],$
where $Q_{(1-\alpha)(1+n^{-1})}(\mu)$ is the $(1-\alpha)(1+n^{-1})$ quantile of the distribution of $\mu$. Possible improvements of this conformity score are also investigated in [22, 34]. In particular, [34] has shown that the constructed interval converges to the narrowest possible bands that achieve conditional coverage under mild assumptions. However, when the conditional distribution of the response has widely separated high-density regions, the ideal prediction set is not necessarily an interval. In such cases, a conformal method should be able to generate disjoint regions instead of being restricted to intervals; see Figure 1 for a simple illustration and [41] for a discussion and examples.
Figure 1: Bimodal example.
This paper is concerned with improving the construction of prediction sets, especially when an estimate of the conditional distribution $\textup{P}_{Y\mid X}$ is available. The approach we propose utilizes the flexibility of conformal prediction methods to generate confidence sets that better capture the structure of the predictive distribution. This leads to improved conditional coverage and efficiency, especially in situations with high heteroskedasticity where the width of the confidence interval can vary significantly between regions.
In order to achieve this, we pursue the fundamental question of how conditionally valid prediction sets can be derived. There are many, mostly negative, results on conditional validity, which can only be achieved under strong assumptions on the joint distribution of $(X,Y)$, which often fail to be satisfied [13]. One possible solution is to divide the space $\mathbb{R}^{d}$ into several regions and learn a specific quantile for each of these regions [15, 16, 1]. However, this approach has significant drawbacks, especially since splitting the space $\mathbb{R}^{d}$ into multiple regions for a given calibration set typically leads to an increase in the length of the prediction set [29, 27]. As shown in [3, 28], the conditional coverage follows a beta distribution. The error deviation is therefore of the order of $1/\sqrt{n_{x}}$, where $n_{x}$ is the number of calibration data contained in the bin associated with $x$. Thus, an accuracy of $0.01$ would require almost $10^{4}$ data points in each bin, which limits the feasibility of binning methods.
Our proposed method aims to overcome these limitations and provide a more feasible and efficient solution for marginally valid prediction sets equipped with some conditional guarantees. In particular, our work addresses these challenges through the following main contributions:
  • We propose a new method for constructing conditional confidence intervals that adapts to the local structure of the data distribution and allows the generation of confidence intervals that are more informative; see Section 2.
@@
  • We develop a theoretical framework to analyze the properties of the proposed method, establishing its approximate conditional validity; see Section 3.
@@
  • We demonstrate the effectiveness of the proposed method through a series of experiments on synthetic and real-world datasets; see Section 4. The results show that it outperforms existing methods in terms of conditional coverage.
2 Conditionally valid Probabilistic Conformal Prediction
Figure 2: Schematic representation of CP2.
Problem Setup and Sketch of the Method.
@@ 2 Conditionally valid Probabilistic Conformal Prediction
Suppose we are given $n$ samples $\{(X_{k},Y_{k})\}_{k=1}^{n}$ and we must now predict the unknown value of $Y_{n+1}$ at a test point $X_{n+1}$. We assume that all samples $\{(X_{k},Y_{k})\}_{k=1}^{n+1}$ are i.i.d. from an arbitrary joint distribution $\textup{P}_{X,Y}$ over the feature vectors $X\in\mathbb{R}^{d}$ and response variables $Y\in\mathcal{Y}$. The target set $\mathcal{Y}$ can be either finite or continuous. Our goal is to construct a prediction set $\mathcal{C}_{\alpha}(X_{n+1})$ that contains the unobserved output $Y_{n+1}$ with probability close to $1-\alpha$, where $\alpha\in(0,1)$ is the user-specified confidence level.
We aim to develop a flexible plug-and-play method that combines existing conformal methods with conditional distribution estimation $\textup{P}_{Y\mid X}$. We want to construct marginally valid predictive sets with approximate conditional validity. There are three main ingredients for our approach:
@@
    We specify a family of confidence sets $\mathcal{R}(x;t)$ parameterized by $t\in\mathbb{R}$. For instance, $t$ can be chosen as the radius of a ball centered around an estimate of conditional mean $\textup{P}_{Y\mid X}$.
@@
    We split the available data into the training and calibration ones. An estimator $\Pi_{Y\mid X}$ of the conditional probability $\textup{P}_{Y\mid X}$ is learnt using the training data; see Remark 2.2. Given $x\in\mathbb{R}^{d}$, if $\Pi_{Y\mid X=x}$ closely approximates the true conditional probability $\textup{P}_{Y\mid X=x}$, then for any $\tau_{x}\in\mathbb{R}$ such that $\Pi_{Y\mid X=x}(\mathcal{R}(x;\tau_{x}))\geq 1-\alpha$, it follows that $\mathcal{R}(x;\tau_{x})$ is an approximately valid prediction set; see Theorem 3.2 for more details.

### v2

# Probabilistic Conformal Prediction with Approximate Conditional Validity

@@ Abstract
We develop a new method for generating prediction sets that combines the flexibility of conformal methods with an estimate of the conditional distribution $\textup{P}_{Y\mid X}$. Existing methods, such as conformalized quantile regression and probabilistic conformal prediction, usually provide only a marginal coverage guarantee. In contrast, our approach extends these frameworks to achieve approximately conditional coverage, which is crucial for many practical applications. Our prediction sets adapt to the behavior of the predictive distribution, making them effective even under high heteroscedasticity. While exact conditional guarantees are infeasible without assumptions on the underlying data distribution, we derive non-asymptotic bounds that depend on the total variation distance of the conditional distribution and its estimate. Using extensive simulations, we show that our method consistently outperforms existing approaches in terms of conditional coverage, leading to more reliable statistical inference in a variety of applications.
@@ 1 Introduction
Conformal predictions are commonly used to construct prediction sets. Under minimal assumptions, they offer finite-sample validity (Vovk et al., 2005; Shafer & Vovk, 2008). However, significant challenges arise with high heteroskedasticity, often leading to incorrect inferences (Dewolf et al., 2023). The split-conformal approach uses a set of $n$ calibration data points $\{(X_{k},Y_{k})\}_{k\in[n]}$ with $X_{k}\in\mathbb{R}^{d}$ and $Y_{k}\in\mathcal{Y}$ to create a prediction set $\mathcal{C}_{\alpha}(x)$ where $\alpha\in(0,1)$. For each $x\in\mathbb{R}^{d}$, the prediction set based on a conformity score function $V\colon\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$, is given by
@@
$\mathcal{C}_{\alpha}(x)=\left\{{y\in\mathcal{Y}\colon V(x,y)\leq Q_{1-\alpha}\left({\frac{1}{n+1}\sum\nolimits_{k=1}^{n}\delta_{V(X_{k},Y_{k})}+\frac{1}{n+1}\delta_{\infty}}\right)}\right\},$ (1)
@@
where $\delta_{x}$ is the Dirac mass and $Q_{1-\alpha}$ is the $(1-\alpha)$-quantile of the adjusted empirical score distribution $\frac{1}{n+1}\sum_{k=1}^{n}\delta_{V(X_{k},Y_{k})}+\frac{1}{n+1}\delta_{\infty}$. If the calibration data $\{(X_{k},Y_{k})\}_{k\in[n]}$ is drawn i.i.d. from a population distribution $\textup{P}_{X,Y}$, then for any new data point $(X_{n+1},Y_{n+1})\sim\textup{P}_{X,Y}$ sampled independently of the calibration data, the conformal theory ensures the marginal validity of $\mathcal{C}_{\alpha}(X_{n+1})$, meaning that $\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)\geq 1-\alpha$. This marginal guarantee can hide significant discrepancies in the coverage of different regions of the input space $\mathbb{R}^{d}$; see e.g. Izbicki et al. (2022). Conditional validity is a more desirable guarantee than marginal validity: for any $x\in\mathbb{R}^{d}$, the set $\mathcal{C}_{\alpha}(x)$ is conditionally valid if
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x}\right)\geq 1-\alpha.$ (2)
However, this property cannot be achieved without further assumptions about the data distribution; see Vovk (2012); Lei & Wasserman (2014). For practical purposes, it is enough to construct sets $\mathcal{C}_{\alpha}$ which approximate (2), and ideally achieve it asymptotically under suitable conditions in the limit of large sample size $n$.
Related work. A great deal of research has been devoted to this problem, starting with the case where $\mathcal{Y}=\mathbb{R}$. For example Romano et al. (2019); Kivaranovic et al. (2020) proposed methods based on estimate of the lower and upper conditional quantile function $\hat{q}_{\alpha/2}$ and $\hat{q}_{1-\alpha/2}$, to define a quantile-based conformity score: $V(x,y)=\max\left\{{\hat{q}_{\alpha/2}(x)-y,y-\hat{q}_{1-\alpha/2}(x)}\right\}$ which is then conformalized. Improvements of this conformity score are investigated in Kivaranovic et al. (2020); Sesia & Candès (2020). Sesia & Candès (2020) have shown that the constructed interval converges to the narrowest possible interval that achieve conditional coverage under mild assumptions. We stress that these methods are specific to the case $\mathcal{Y}=\mathbb{R}$; in addition, when the conditional distribution $\textup{P}_{Y\mid X}$ is multimodal, restricting prediction to intervals is suboptimal; see Wang et al. (2023) for a discussion and examples. It has been suggested in Guan (2023); Alaa et al. (2023) to partition $\mathbb{R}^{d}$ and learn a specific quantile for each regions. Splitting the space $\mathbb{R}^{d}$ into multiple regions typically leads to an increase in the length of the prediction set; see Romano et al. (2020a); Melki et al. (2023) for comments.
A number of studies have focused on the construction of prediction sets using an estimator of the conditional distribution $\textup{P}_{Y\mid X}$. Again, in the case where $\mathcal{Y}=\mathbb{R}$, Cai et al. (2014); Lei & Wasserman (2014) have constructed prediction intervals based on an estimator of the conditional density and have established the asymptotic validity under appropriate conditions. Han et al. (2022) use kernel density estimation to construct asymmetric prediction bands. However, this method cannot handle bimodality as it generates a single interval. On the other hand, Sesia & Romano (2021) partition the domain of $Y$ into bins to create a histogram approximation of $\textup{P}_{Y\mid X}$. The authors showed that their method satisfies the marginal validity while achieving the asymptotic conditional coverage; see also Lei et al. (2018). Asymptotic conditional coverage is also obtained in Sesia & Candès (2020); Cauchois et al. (2020) using quantile regression-based methods, or using cumulative distribution function estimators (Izbicki et al., 2020; Chernozhukov et al., 2021). Conditionally valid prediction sets have been shown to improve the robustness to perturbations (Gendler et al., 2021). Guha et al. (2024) proposed a novel approach that converts regression tasks into classification problems by binning the output space and discretizing the labels. Leveraging this discretization, they approximate the conditional density to construct prediction sets that correspond to regions of Highest Predictive Density (HPD). A key limitation of these methods lies in the discretization process, as the number of labels required for complex scenarios can become computationally prohibitive. Diamant et al. (2024) introduced an approach that estimates conditional densities using neural networks parameterized by splines, offering a more flexible representation.
Very few studies have addressed the scenario where the prediction target is multi-dimensional, i.e., $\mathcal{Y}=\mathbb{R}^{q}$ with $q>1$. Wang et al. (2023) developed the PCP method, based on implicit conditional generative models (CGMs). These CGMs allow for the generation of samples from the conditional distribution without requiring an explicit closed-form expression. The PCP method constructs prediction sets as unions of balls, whose centers are generated from the CGM. However, in PCP, the radius of these balls is fixed across the space, which can introduce significant limitations. In regions with low variability, a fixed radius may result in over-coverage, while for highly dispersed conditional distributions, it may lead to under-coverage, failing to capture the full extent of the relevant space. We observed that the performance of PCP deteriorates as the number of balls increases, exacerbating the heteroscedasticity problem. This underscores the need for a more adaptive methodology that can dynamically adjust the size of prediction sets in response to local variability in the data.
Our work addresses these challenges through the following main contributions:
  • We propose a new CP2 method for constructing conditional confidence sets that adapts to the local structure of the data distribution, capable of addressing both classical regression problems where $\mathcal{Y}=\mathbb{R}$ and more complex multi-dimensional prediction tasks where $\mathcal{Y}=\mathbb{R}^{q}$. Our approach is versatile in accommodating scenarios involving either an explicit conditional density estimator or an implicit generative model; see Section 2.
@@
  • We develop a theoretical framework to analyze the properties of the proposed CP2 method, establishing both its marginal and approximate conditional validity. Furthermore, we demonstrate that asymptotic conditional coverage is attainable under a weak consistency assumption on the predictive distribution; see Section 3.
@@
  • We demonstrate the effectiveness of the proposed method through a series of experiments on synthetic and real-world datasets. The results indicate that our approach consistently outperforms existing methods in terms of conditional coverage. Specifically, it excels in handling classical regression problems, effectively addressing multimodality, and proves robust in the more challenging setting of multidimensional prediction tasks; see Section 4.
2 The CP2 framework
@@
We want to construct marginally valid predictive sets with approximate conditional validity. We follow the split-conformal approach to conformal inference due to its computational feasibility with large datasets; see along others Papadopoulos et al. (2002); Papadopoulos (2008); Papadopoulos et al. (2011); Romano et al. (2019); Kivaranovic et al. (2020). The first step involves splitting the data samples into two disjoint subsets, the training set $\mathcal{T}=\{(\tilde{X}_{k},\tilde{Y}_{k})\}_{k=1}^{m}$ and calibration set $\mathcal{C}=\{(X_{k},Y_{k})\}_{k=1}^{n}$. It is assumed in the sequel that the training and calibration data are mutually independent and i.i.d. with distribution $\textup{P}_{X,Y}$ over the feature vectors $X\in\mathbb{R}^{d}$ and response variables $Y\in\mathcal{Y}$. The target set $\mathcal{Y}$ can be either finite or continuous. We must now predict the unknown value of $Y_{n+1}$ given an input $X_{n+1}$, independent from $\mathcal{T}\cup\mathcal{C}$. Our goal is to construct a prediction set $\mathcal{C}_{\alpha}(X_{n+1})$ that contains the unobserved output $Y_{n+1}$ with probability close to $1-\alpha$, where $\alpha\in(0,1)$ is the user-specified confidence level.
An estimator $\Pi_{Y\mid X}$ of the conditional probability $\textup{P}_{Y\mid X}$ is learnt using the training data. There is a rich body of research on nonparametric conditional density estimation, with the most common methods relying on smoothing techniques such as kernel smoothing and local polynomial fitting. An alternative approach involves transforming the conditional density estimation task into a regression problem, allowing the application of nonparametric regression methods to approximate the conditional density. More recently, generative methods leveraging deep neural networks have been developed for nonparametric conditional density estimation, which enable sampling from the conditional distribution; see Abadi et al. (2016); Zhou et al. (2021) for examples of these techniques. In the sequel, the choice of this algorithm is treated as a black box. There are three main ingredients for our approach:
@@
    In classical conformal prediction methods, the shape of the prediction set is specified by the score function $V(x,y)$; see (1). The first element of our construction is a family of explicitly given confidence sets $\mathcal{R}_{z}(x;t)$ parameterized by $t\in\mathsf{T}$ where $\mathsf{T}$ is a subset of $\mathbb{R}$ and $z\in\mathcal{Z}$ auxiliary variables. The index set $\mathsf{T}$ can, in most cases, be taken as either $\mathsf{T}=\mathbb{R}$ or $\mathsf{T}=\mathbb{R}_{+}$. The key assumptions for the confidence set are as follows: (a) The size of $\mathcal{R}_{z}(x;t)$ increases with $t\in\mathsf{T}$ for any $z\in\mathcal{Z}$. In addition, by choosing a sufficiently large value of $t$, the entire output space $\mathcal{Y}$ can be covered. (b) There exists a form of continuity for $t\mapsto\mathcal{R}_{z}(x;t)$. In mathematical terms, the following assumption should hold:
    H 1.
    For any $(x,z)\in\mathbb{R}^{d}\times\mathcal{Z}$, the confidence sets $\{\mathcal{R}_{z}(x;t)\}_{t\in\mathsf{T}}$ are non-decreasing, $\Pi_{Y|X=x}(\cap_{t\in\mathsf{T}}\mathcal{R}_{z}(x;t))=0$, $\cup_{t\in\mathcal{T}}\mathcal{R}_{z}(x;t)=\mathcal{Y}$; in addition, for any $t\in\mathsf{T}$, $\cap_{t^{\prime}>t}\mathcal{R}_{z}(x;t^{\prime})=\mathcal{R}_{z}(x;t)$.
    Example 2.1.
    For instance, $\mathcal{R}(x;t)$ can be chosen as a ball centered around an estimate of conditional mean $\textup{P}_{Y\mid X}$ with radius $t$ and there is no auxiliary variables then (we then remove subscript $z$). Examples of confidence intervals specialized to the case where $\mathcal{Y}=\mathbb{R}$ are given in Table 1.
    Example 2.2.
    If the predictive distribution is multimodal, a ball centered around the predictive mean often fails to provide an informative prediction set. Ideally, $\mathcal{R}_{z}(x;t)$ should correspond to the set with the highest predictive density (HPD) of $\textup{P}_{Y\mid X}$. However, HPD regions are difficult to determine in practice, even when the conditional predictive density is available. Following Wang et al. (2023), we may define the prediction sets as $\mathcal{R}_{z}(x;t):=\cup_{i=1}^{M}\mathrm{B}(y_{i},t)$, and they depends on an exogenous variables $z=(y_{1},\ldots,y_{M})\in\mathcal{Z}$, where each $y_{i}$ is sampled conditionally independently from the conditional generative model $\Pi_{Y\mid X=x}$.
    In a general setting, the auxiliary variables $z\in\mathcal{Z}$ are sampled according to a kernel $\bar{\Pi}_{Z\mid X=x}$ for a given $x$.
@@
    To localize conformal prediction methods, it is convenient to introduce a function $f_{\tau}(\lambda)$ parameterized by $\tau$. Such function aims to transform the conformity score $\lambda$ and was introduced (albeit in a slightly different form) in Deutschmann et al. (2023); Han et al. (2022). Examples of such a function are $f_{\tau}(\lambda)=\tau\lambda$ and $f_{\tau}(\lambda)=\tau+\lambda$. We assume that:
    H 2.
    There exists $\varphi\in\mathsf{T}$ such that $\tau\in\mathsf{T}\mapsto f_{\tau}(\varphi)$ is increasing and bijective. In addition, $\lambda\in\mathsf{T}\mapsto f_{\tau}(\lambda)$ is increasing for any $\tau\in\mathsf{T}$.
    We define $\tau_{x,z}$ using the estimated predictive density $\Pi_{Y\mid X=x}$ according to
    $\tau_{x,z}=\inf\left\{{\tau\in\mathsf{T}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha}\right\}.$ (3)
    It is easily shown that for any $\alpha\in(0,1)$, $x\in\mathbb{R}^{d}$, $z\in\mathcal{Z}$, then $\tau_{x,z}\in\mathsf{T}$ and $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))\geq 1-\alpha$; see Lemma A.3.